OpenStudy (anonymous):

The average of two numbers, a and b, is given by: A=a+b/2 If the average is 93, and one number is 88, what is the other number? Please show work

OpenStudy (anonymous):

solve \[93=\frac{88+x}{2}\] \[88+x=2\times 93=186\] \[x=186-88=98\]